# Timothy M. Chan UIUC

### v2

@@ 1 Introduction
However, so far not even poly-logarithmic improvements of Abrahamson’s algorithm have been reported. This is not due to a lack of interest. In fact, many algorithms are designed to shave logarithmic factors for stringology problems (e.g. [${}^{+}$20, MP80, Mye92, Ind98, CLZ03, BF08, BT09, Gra16]). In this paper, we will tackle the following decades-old question:
@@
More recently, Chan, Golan, Kociumaka, Kopelowitz and Porat in STOC’20 [${}^{+}$20] partially answered 2: when the pattern length $m$ satisfies $m\geq\varepsilon^{-28}$, one can $(1+\varepsilon)$-approximate Text-to-Pattern Hamming Distances in $\widetilde{O}(n)$ time, without any $\varepsilon^{-O(1)}$ factors. The assumption may be relaxed to $m\geq\varepsilon^{-10}$ if the matrix multiplication exponent $\omega$ is equal to 2, and if the goal is to obtain better than $\widetilde{O}(\varepsilon^{-1}n)$ time instead of $\widetilde{O}(n)$, the assumption can be relaxed further by re-analyzing/modifying their algorithm. However, inherently their approach is unable to beat $\varepsilon^{-1}n$ if $\varepsilon^{-1}$ is large, for example, when $\varepsilon^{-1}$ is $m^{1/3}$ or $\sqrt{m}$. The $\varepsilon^{-1}=\sqrt{m}$ case is particularly instructive: here, $\widetilde{O}(\varepsilon^{-1}n)$ coincides with the $\widetilde{O}(n\sqrt{m})$ bound for the exact problem; for distances that are $\Theta(m)$, we are demanding $O(\sqrt{m})$ additive error, and sampling-based approaches do not seem to offer any speedup (if we try to estimate distances by sampling different positions of the pattern string, we would need a sample size of $\Omega(m)$, which is not any smaller than the length of the original pattern string).
Chan et al. [${}^{+}$20] also gave an $O(\varepsilon^{-2}n)$ time randomized algorithm (correct with high probability) without any logarithmic factors, which is preferable when $\varepsilon^{-1}$ is small. Both 1 and 2 were explicitly asked during a talk on [${}^{+}$20] given by Kociumaka. [^5] Talk video link: https://youtu.be/WEiQjjTBX-4?t=2820
Other variations of $(1+\varepsilon)$-approximation text-to-pattern matching problem have also been studied in the literature, such as replacing Hamming distance by other $\ell_{p}$ norms [LP11, GU18, SU19, Uzn20a], or restricting to algorithms that do not use FFT [${}^{+}$20, Uzn20a]. See also the survey by Uznański [Uzn20b].
@@
After a long line of works [LV86, LV89, GG86, SV96, CH02, ALP04, ${}^{+}$16, GU18, ${}^{+}$20] (see Table 1), the current fastest algorithm by Chan, Golan, Kociumaka, Kopelowitz and Porat [${}^{+}$20] is a Monte Carlo randomized algorithm (correct with high probability) in $O\big{(}n+\min\big{(}\frac{nk}{\sqrt{m}}\sqrt{\log m},\frac{nk^{2}}{m}\big{)}\big{)}$ time, shaving off some logarithmic factors from the earlier deterministic algorithm by Gawrychowski and Uznański [GU18] in $O\big{(}n\log^{2}m\log\sigma+\frac{nk\sqrt{\log n}}{\sqrt{m}}\big{)}$ time. Gawrychowski and Uznański [GU18] also extended Indyk’s fine-grained reduction (mentioned in the notes of Clifford [Cli09]) to show a tight conditional lower bound for combinatorial algorithms solving the $k$-mismatch problem under the Boolean Matrix Multiplication Hypothesis.
@@ 1.1 Our results
This speeds up the previous $O\big{(}n+\min\big{(}\frac{nk}{\sqrt{m}}\sqrt{\log m},\frac{nk^{2}}{m}\big{)}\big{)}$-time Monte Carlo algorithm (with high success probability) time by Chan, Golan, Kociumaka, Kopelowitz and Porat [${}^{+}$20], and cleans up all the extra factors from the long line of previous works shown in Table 1 (note that $n+\frac{nk^{2}}{m}$ is never better than $n+\frac{nk}{\sqrt{m}}$).
@@
Clifford, Fontaine, Porat, Sach, and Starikovskaya [${}^{+}$16] $O(n\log^{O(1)}m+\frac{nk^{2}\log k}{m})$
@@
Chan, Golan, Kociumaka, Kopelowitz and Porat [${}^{+}$20] $O(n+\min\{\frac{nk\sqrt{\log m}}{\sqrt{m}},\,\frac{nk^{2}}{m}\})$
@@
Chan, Golan, Kociumaka, Kopelowitz and Porat [${}^{+}$20] $\widetilde{O}(n)$ for $m\gg\varepsilon^{-28}$ random sampling $+$ rect. matrix mult.
@@ 1.2 Technical overview
Our new approximation algorithm for Theorem 1.1 uses an approach that is markedly different from all previous approximation algorithms. The algorithms by Karloff [Kar93] and Kopelowitz and Porat [KP18] used random projection to reduce the alphabet size; afterwards, the problem can be solved by FFT. Karloff’s algorithm required $O(\varepsilon^{-2})$ projections, whereas Kopelowitz and Porat’s algorithm required a reduced alphabet size of $O(\varepsilon^{-1})$. On the other hand, the algorithms by Indyk [Ind98] and Chan et al. [${}^{+}$20] used random sampling to examine selected positions of the pattern and text strings. An application of the Chernoff bound leads to $O(\varepsilon^{-2})$ factors in the running time, which are too big for the critical case $\varepsilon^{-1}=\sqrt{m}$.
@@
    Its use of matrix multiplication is non-trivial and interesting. It is open whether fast matrix multiplication helps for the exact Text-to-Pattern Hamming Distances problem, but our new algorithm demonstrates that it helps for the approximate problem. Chan et al.’s previous algorithm [${}^{+}$20] also used rectangular matrix multiplication to speed up certain steps, but our algorithm here relies on matrix multiplication (via Equality Product) in a more essential way.
@@
It is straightforward to combine our new lemma with existing algorithms [GU18, ${}^{+}$20] to obtain our result on the $k$-mismatch problem (Theorem 1.5) and the Dominance Convolution problem (Theorem 1.4).
@@ 3.4 Approximate Counting Colored All-Numbers $3$SUM
[…]
@@ 3.5 Applying to Text-to-Pattern Hamming Distances
For all shifts with Hamming distance $k\leq\varepsilon^{-\gamma}\sqrt{m}$, we can use a known exact algorithm running in $\widetilde{O}(\frac{n}{m}\cdot(m+k\sqrt{m}))\leq\widetilde{O}(\varepsilon^{-\gamma}n)$ time [GU18, ${}^{+}$20]. Otherwise, the bound in Theorem 3.6 is at most $\widetilde{O}(\frac{n}{m}\cdot(\varepsilon^{-1+\beta}m+\varepsilon^{-1-\beta+2\beta\gamma}m))$. We thus obtain an upper bound of $\widetilde{O}(\varepsilon^{-\gamma}n)$ in all cases, by setting $\gamma:=(1+\beta)/(1+2\beta)=8/(11-\omega)<0.928$. We run the entire algorithm for every $k$ that is a power of 2. ∎
@@ 4.1 $X+Y$ Lemma
[…]
@@ 4.3 $k$-Mismatch
Our technique also improves the previous algorithm for the $k$-mismatch problem. In fact, we only need to replace the use of FFT in [${}^{+}$20]’s $O\big{(}n+\min\big{(}\frac{nk}{\sqrt{m}}\sqrt{\log m},\frac{nk^{2}}{m}\big{)}\big{)}$-time algorithm with our new Lemma 4.1.
@@
The bottleneck of [${}^{+}$20]’s algorithm lies in the following task: given $2t$ sparse sequences $f_{1},\ldots,f_{t},g_{1},\ldots,g_{t}$ whose supports are all in $[n]$, and the total size of their supports is $O(k)$, compute (a sparse representation of) $f_{i}\star g_{i}$ for every $i$. Additionally, all nonzero entries of $f_{i}$ and $g_{i}$ are either 00 or $1$. The running time for this task in [${}^{+}$20, Lemma 7.8] is $O(k\min(k,\sqrt{n\log n}))$. It suffices to improve it to provide an $O(k\min(k,\sqrt{n}))$.
@@ 5 Deterministic Exact Text-to-Pattern Hamming Distances
Given three integer sequences $A,B,C$ of length $U$, with the promise that $(A\star B-C)[i]\geq 0$ for every $i$, we can compute $A\star B$ in
@@
deterministic time, where $t=\max\{||A||_{0},||B||_{0},||A\star B-C||_{0}\}$.
@@
Without loss of generality, assume $U$ is a power of $2$. If $U=1$, then the problem can be trivially solved in $O(1)$ time. In the following, we assume $U\geq 2$.
Let $U^{\prime}=U/2$. For every $i\in[U^{\prime}]$, let $A^{\prime}[i]:=A[i]+A[i+U^{\prime}]$. We similarly prepare $B^{\prime}$ and $C^{\prime}$. Then we recursively compute $A^{\prime}\star B^{\prime}$.
Given $A^{\prime}\star B^{\prime}$, we can compute $A^{\prime}\star B^{\prime}-C^{\prime}$ in $O(U)$ time. Let $S$ be the support of $A^{\prime}\star B^{\prime}-C^{\prime}$. Notice that $(A^{\prime}\star B^{\prime}-C^{\prime})[i]=(A\star B-C)[i]+(A\star B-C)[i+U^{\prime}]+(A\star B-C)[i+2U^{\prime}]+(A\star B-C)[i+3U^{\prime}]$. Therefore, by setting $T:=\bigcup_{s\in S}\{s,s+U^{\prime},s+2U^{\prime},s+3U^{\prime}\}$, $T$ is guaranteed to be a superset of the support of $A\star B-C$. Furthermore, $|T|=O(||A\star B-C||_{0})=O(t)$.
Next, we apply Lemma 5.1 on $T$ and $U$ and find $r=2^{O(\sqrt{\log t\log\log U})}$ integers $m_{1},\ldots,m_{r}=t\cdot 2^{\Theta(\sqrt{\log t\log\log U})}\cdot\operatorname{\mathrm{poly}}\log(t,U)$. For each $k\in[r]$, we prepare two arrays $A_{k}$ and $B_{k}$, which are defined as $A_{k}[i]:=\sum_{j\equiv i\bmod{m_{k}}}A[j]$ and $B_{k}[i]:=\sum_{j\equiv i\bmod{m_{k}}}B[j]$. Then we compute the following array $D_{k}$ via FFT in $\widetilde{O}(m_{k})$ time:
@@
The recursion adds an $\log(U)$ factor to the $t\cdot 2^{O(\sqrt{\log t\log\log U})}\cdot\operatorname{\mathrm{poly}}\log(t,U)$ part of the running time. It only adds a constant factor to the $O(U)$ part of the running time, as $U$ is halved at each recursion level. ∎
@@
Let $p$ be a prime whose range is to be determined later. First, we use Lemma 2.2 to compute $\mbox{\sc count}[z]\bmod{p}$ for every $z\in[U]$ in $O(U\log p)$ time.
Then we apply Lemma 5.2, using $X$ for $A$, $Y$ for $B$ and $\mbox{\sc count}[z]\bmod{p}$ for $C$. Clearly, $||A\star B-C||_{0}\leq\frac{n^{2}}{p}$, so the overall running time for computing $A\star B$ (which gives $\mbox{\sc count}[z]$) is $O(U)+t\cdot 2^{O(\sqrt{\log t\log\log U})}\cdot\operatorname{\mathrm{poly}}\log(t,U)$ for $t=\max\{n,\frac{n^{2}}{p}\}$. Also, as $2U<n^{2}$, we can simplify the running time to $O\left(U+t\cdot 2^{c\sqrt{\log t\log\log U})}\right)$ for some constant $c$.
Finally, picking $p$ from $\Theta\left(\frac{n^{2}}{U}\cdot 2^{c\sqrt{\log n\log\log U}}\right)$ (it only takes $n^{o(1)}$ time to find such a prime as $\frac{n^{2}}{U}=n^{o(1)}$) gives the desired running time. ∎

### v3

@@ 1 Introduction
However, so far not even poly-logarithmic improvements of Abrahamson’s algorithm have been reported. This is not due to a lack of interest. In fact, many algorithms are designed to shave logarithmic factors for stringology problems (e.g. [CGK+20, MP80, Mye92, Ind98, CLZ03, BF08, BT09, Gra16]). In this paper, we will tackle the following decades-old question:
@@
More recently, Chan, Golan, Kociumaka, Kopelowitz and Porat in STOC’20 [CGK+20] partially answered 2: when the pattern length $m$ satisfies $m\geq\varepsilon^{-28}$, one can $(1+\varepsilon)$-approximate Text-to-Pattern Hamming Distances in $\widetilde{O}(n)$ time, without any $\varepsilon^{-O(1)}$ factors. The assumption may be relaxed to $m\geq\varepsilon^{-10}$ if the matrix multiplication exponent $\omega$ is equal to 2, and if the goal is to obtain better than $\widetilde{O}(\varepsilon^{-1}n)$ time instead of $\widetilde{O}(n)$, the assumption can be relaxed further by re-analyzing/modifying their algorithm. However, inherently their approach is unable to beat $\varepsilon^{-1}n$ if $\varepsilon^{-1}$ is large, for example, when $\varepsilon^{-1}$ is $m^{1/3}$ or $\sqrt{m}$. The $\varepsilon^{-1}=\sqrt{m}$ case is particularly instructive: here, $\widetilde{O}(\varepsilon^{-1}n)$ coincides with the $\widetilde{O}(n\sqrt{m})$ bound for the exact problem; for distances that are $\Theta(m)$, we are demanding $O(\sqrt{m})$ additive error, and sampling-based approaches do not seem to offer any speedup (if we try to estimate distances by sampling different positions of the pattern string, we would need a sample size of $\Omega(m)$, which is not any smaller than the length of the original pattern string).
Chan et al. [CGK+20] also gave an $O(\varepsilon^{-2}n)$ time randomized algorithm (correct with high probability) without any logarithmic factors, which is preferable when $\varepsilon^{-1}$ is small. Both 1 and 2 were explicitly asked during a talk on [CGK+20] given by Kociumaka. [^5] Talk video link: https://youtu.be/WEiQjjTBX-4?t=2820
Other variations of $(1+\varepsilon)$-approximation text-to-pattern matching problem have also been studied in the literature, such as replacing Hamming distance by other $\ell_{p}$ norms [LP11, GU18, SU19, Uzn20a], or restricting to algorithms that do not use FFT [CGK+20, Uzn20a]. See also the survey by Uznański [Uzn20b].
@@
After a long line of works [LV86, LV89, GG86, SV96, CH02, ALP04, CFP+16, GU18, CGK+20] (see Table 1), the current fastest algorithm by Chan, Golan, Kociumaka, Kopelowitz and Porat [CGK+20] is a Monte Carlo randomized algorithm (correct with high probability) in $O\big{(}n+\min\big{(}\frac{nk}{\sqrt{m}}\sqrt{\log m},\frac{nk^{2}}{m}\big{)}\big{)}$ time, shaving off some logarithmic factors from the earlier deterministic algorithm by Gawrychowski and Uznański [GU18] in $O\big{(}n\log^{2}m\log\sigma+\frac{nk\sqrt{\log n}}{\sqrt{m}}\big{)}$ time. Gawrychowski and Uznański [GU18] also extended Indyk’s fine-grained reduction (mentioned in the notes of Clifford [Cli09]) to show a tight conditional lower bound for combinatorial algorithms solving the $k$-mismatch problem under the Boolean Matrix Multiplication Hypothesis.
@@ 1.1 Our results
This speeds up the previous $O\big{(}n+\min\big{(}\frac{nk}{\sqrt{m}}\sqrt{\log m},\frac{nk^{2}}{m}\big{)}\big{)}$-time Monte Carlo algorithm (with high success probability) time by Chan, Golan, Kociumaka, Kopelowitz and Porat [CGK+20], and cleans up all the extra factors from the long line of previous works shown in Table 1 (note that $n+\frac{nk^{2}}{m}$ is never better than $n+\frac{nk}{\sqrt{m}}$).
@@
Clifford, Fontaine, Porat, Sach, and Starikovskaya [CFP+16] $O(n\log^{O(1)}m+\frac{nk^{2}\log k}{m})$
@@
Chan, Golan, Kociumaka, Kopelowitz and Porat [CGK+20] $O(n+\min\{\frac{nk\sqrt{\log m}}{\sqrt{m}},\,\frac{nk^{2}}{m}\})$
@@
Chan, Golan, Kociumaka, Kopelowitz and Porat [CGK+20] $\widetilde{O}(n)$ for $m\gg\varepsilon^{-28}$ random sampling $+$ rect. matrix mult.
@@ 1.2 Technical overview
Our new approximation algorithm for Theorem 1.1 uses an approach that is markedly different from all previous approximation algorithms. The algorithms by Karloff [Kar93] and Kopelowitz and Porat [KP18] used random projection to reduce the alphabet size; afterwards, the problem can be solved by FFT. Karloff’s algorithm required $O(\varepsilon^{-2})$ projections, whereas Kopelowitz and Porat’s algorithm required a reduced alphabet size of $O(\varepsilon^{-1})$. On the other hand, the algorithms by Indyk [Ind98] and Chan et al. [CGK+20] used random sampling to examine selected positions of the pattern and text strings. An application of the Chernoff bound leads to $O(\varepsilon^{-2})$ factors in the running time, which are too big for the critical case $\varepsilon^{-1}=\sqrt{m}$.
@@
    Its use of matrix multiplication is non-trivial and interesting. It is open whether fast matrix multiplication helps for the exact Text-to-Pattern Hamming Distances problem, but our new algorithm demonstrates that it helps for the approximate problem. Chan et al.’s previous algorithm [CGK+20] also used rectangular matrix multiplication to speed up certain steps, but our algorithm here relies on matrix multiplication (via Equality Product) in a more essential way.
@@
It is straightforward to combine our new lemma with existing algorithms [GU18, CGK+20] to obtain our result on the $k$-mismatch problem (Theorem 1.5) and the Dominance Convolution problem (Theorem 1.4).
@@ 3.4 Approximate Counting Colored All-Numbers $3$SUM
$\widetilde{O}\left(\varepsilon^{-1+\beta}n^{1-\beta}k^{\beta}+\varepsilon^{-1-\beta}n^{1+\beta}/k^{2\beta}+n\right),\qquad\mbox{where $\beta:=(3-\omega)/(5+\omega)$.}$
@@ 3.5 Applying to Text-to-Pattern Hamming Distances
For all shifts with Hamming distance $k\leq\varepsilon^{-\gamma}\sqrt{m}$, we can use a known exact algorithm running in $\widetilde{O}(\frac{n}{m}\cdot(m+k\sqrt{m}))\leq\widetilde{O}(\varepsilon^{-\gamma}n)$ time [GU18, CGK+20]. Otherwise, the bound in Theorem 3.6 is at most $\widetilde{O}(\frac{n}{m}\cdot(\varepsilon^{-1+\beta}m+\varepsilon^{-1-\beta+2\beta\gamma}m))$. We thus obtain an upper bound of $\widetilde{O}(\varepsilon^{-\gamma}n)$ in all cases, by setting $\gamma:=(1+\beta)/(1+2\beta)=8/(11-\omega)<0.928$. We run the entire algorithm for every $k$ that is a power of 2. ∎
@@ 4.1 $X+Y$ Lemma
    There is a fixed set $\Delta_{h}$ of $O(1)$ size, and mappings $\tau_{h}\colon[U]\rightarrow[q^{O(1)}]$ and $\phi_{h}\colon[q^{O(1)}]\times[q^{O(1)}]\rightarrow\Delta_{h}\cup\{\mbox{undefined}\}$ such that for all $x,y$, $h(x)+h(y)-h(x+y)=\phi_{h}(\tau_{h}(x),\tau_{h}(y))$ if $\phi_{h}(\tau_{h}(x),\tau_{h}(y))$ is defined. Call $(x,y)$ good if $\phi_{h}(\tau_{h}(x),\tau_{h}(y))$ is defined, and bad otherwise.
@@ 4.3 $k$-Mismatch
Our technique also improves the previous algorithm for the $k$-mismatch problem. In fact, we only need to replace the use of FFT in [CGK+20]’s $O\big{(}n+\min\big{(}\frac{nk}{\sqrt{m}}\sqrt{\log m},\frac{nk^{2}}{m}\big{)}\big{)}$-time algorithm with our new Lemma 4.1.
@@
The bottleneck of [CGK+20]’s algorithm lies in the following task: given $2t$ sparse sequences $f_{1},\ldots,f_{t},g_{1},\ldots,g_{t}$ whose supports are all in $[n]$, and the total size of their supports is $O(k)$, compute (a sparse representation of) $f_{i}\star g_{i}$ for every $i$. Additionally, all nonzero entries of $f_{i}$ and $g_{i}$ are either 00 or $1$. The running time for this task in [CGK+20, Lemma 7.8] is $O(k\min(k,\sqrt{n\log n}))$. It suffices to improve it to provide an $O(k\min(k,\sqrt{n}))$.
@@ 5 Deterministic Exact Text-to-Pattern Hamming Distances
Given three integer sequences $A,B,C$ of length $U$ and a set $T\subseteq[U]$, with the promise that $(A\star B-C)[i]\geq 0$ for every $i$ and $T$ is a superset of the support of $A\star B-C$, we can compute $A\star B$ in
@@
deterministic time, where $t=\max\{||A||_{0},||B||_{0},||T||_{0}\}$.
@@
We apply Lemma 5.1 on $T$ and $U$ and find $r=2^{O(\sqrt{\log t\log\log U})}$ integers $m_{1},\ldots,m_{r}=t\cdot 2^{\Theta(\sqrt{\log t\log\log U})}\cdot\operatorname{\mathrm{poly}}\log(t,U)$. For each $k\in[r]$, we prepare two arrays $A_{k}$ and $B_{k}$, which are defined as $A_{k}[i]:=\sum_{j\equiv i\bmod{m_{k}}}A[j]$ and $B_{k}[i]:=\sum_{j\equiv i\bmod{m_{k}}}B[j]$. Then we compute the following array $D_{k}$ via FFT in $\widetilde{O}(m_{k})$ time:
@@
Let $p$ be a prime whose range is to be determined later. We use Lemma 2.2 to compute $\mbox{\sc count}[z]\bmod{p}$ for every $z\in[U]$ in $O(U\log p)$ time.
Then we aim to apply Lemma 5.2 using $X$ for $A$, $Y$ for $B$ and $\mbox{\sc count}[z]\bmod{p}$ for $C$. However, we need a small set $T$ that is a superset of the support of $A\star B-C$. To do so, we recursively solve the following problem.
Let $U^{\prime}=U/2$, and define $X^{\prime}:=\{x\bmod{\frac{U^{\prime}}{2}}:x\in X\}$ and $Y^{\prime}:=\{y\bmod{\frac{U^{\prime}}{2}}:y\in Y\}$. Then we compute $\mbox{\sc count}^{\prime}[z]:=|\{(x,y)\in X^{\prime}\times Y^{\prime}:x+y=z\}|$ recursively in $Q(n,U/2)$ time, where $Q(n,U)$ denotes the running time to solve the instance as described in the lemma statement. We let $T$ be $\{z\in U^{\prime}:\mbox{\sc count}^{\prime}[z]\geq p/2\}+\{0,U^{\prime}/2,U^{\prime},3U^{\prime}/2\}$.
We verify that $T$ is a superset of the support of $A\star B-C$. Note that if $(A\star B-C)[i]\neq 0$ for some $i$, then $\mbox{\sc count}[i]\geq p$, or equivalently, |$|\{(x,y)\in X\times Y:x+y=i\}|\geq p$. This implies
|$\ |\{(x^{\prime},y^{\prime})\in X^{\prime}\times Y^{\prime}:x^{\prime}+y^{\prime}\equiv i\pmod{U^{\prime}/2}\}|$
$=$ |$\ \left|\left\{(x,y)\in X\times Y:\left(x\bmod\frac{U^{\prime}}{2}\right)+\left(y\bmod\frac{U^{\prime}}{2}\right)\equiv i\pmod{U^{\prime}/2}\right\}\right|$
$\geq$ $\ p.$
Therefore, $\mbox{\sc count}^{\prime}[i\bmod{\frac{U^{\prime}}{2}}]\geq p/2$ or $\mbox{\sc count}^{\prime}[(i\bmod{\frac{U^{\prime}}{2}})+(U^{\prime}/2)]\geq p/2$. In either case, $i$ will be included in $T$.
Additionally, it is easy to see that $|T|=O(n^{2}/p)$, as $\sum_{z}\mbox{\sc count}^{\prime}[z]\leq n^{2}$. Then we apply Lemma 5.2, using $X$ for $A$, $Y$ for $B$, $\mbox{\sc count}[z]\bmod{p}$ for $C$ and $T$. The running time is $O(U)+t\cdot 2^{O(\sqrt{\log t\log\log U})}\cdot\operatorname{\mathrm{poly}}\log(t,U)$ for $t=\max\{n,\frac{n^{2}}{p}\}$. Also, as $2U<n^{2}$, we can simplify the running time to $O\left(U+t\cdot 2^{c\sqrt{\log t\log\log U})}\right)$ for some constant $c$.
The overall running time is thus
$Q(n,U/2)+O\left(U\log p+t\cdot 2^{c\sqrt{\log t\log\log U})}\right)$
for $t=\max\{n,\frac{n^{2}}{p}\}$. Picking $p$ from $\Theta\left(\frac{n^{2}}{U}\cdot 2^{c\sqrt{\log n\log\log U}}\right)$ (it only takes $n^{o(1)}$ time to find such a prime as $\frac{n^{2}}{U}=n^{o(1)}$) simplifies the running time to
$Q(n,U/2)+O\left(U\log(n^{2}/U)+U\sqrt{\log n\log\log U}\right),$
and it is easy to see that the recursion only incurs a constant factor blowup to the running time. ∎